Baddy1:

An angle measuring 60 degrees is measured as 62 degrees. What is the percentage error correct to 3 significant figures? A.35.5% B.3.33% C.3.23% D.37.2%

Baddy1:

hope i can get a correct answer

xxemilianaxx:

A. 35.5% :https://www.calculator.net/percent-error-calculator.html

xxemilianaxx:

I thought it would be 35.5?

mhanifa:

It should be (62 - 60)/60

Baddy1:

@extrinix wrote:
So we need to take 2 out of the 62, so we would just divide. \(\dfrac{2}{62}\) Which would give us \(0.03225\). Then we need to move it over 2 decimal places. So we would end up with \(3.225\), which when converted to 3 significant figures, we get \(3.23\), which is C, so C would be your answer.
thank you

Baddy1:

@extrinix wrote:
Difference of the two, which is 2.
thank you

Baddy1:

@xxemilianaxx wrote:
I thought it would be 35.5?
thank you

mhanifa:

2/60= 0.0333 3.33% is correct one

Baddy1:

@mhanifa wrote:
2/60= 0.0333 3.33% is correct one
thank you

mhanifa:

You are welcome!